[Hallc_running] [Revised Logentry] RC Daily Update

rmontgom at jlab.org rmontgom at jlab.org
Sat Nov 24 11:05:02 EST 2018

Logentry Text:
- Swing yesterday completed production data taking on Group 12 (Q2 = 5.0, x = 0.5, pi+).
- SHMS was moved for electron singles data collection. During rotation of spectrometers there was a drive controller error. Expert advice was sought (S. Lassiter) and an escorted access was made to reset the shunt trip breakers, which worked. This was followed by a ~3h30 downtime due to NL13 issues from ~21:00.
- Owl crew completed electron singles data and moved to Group 8 (Q2 = 3.979, x = 0.5, pi-). At  ~5am there was a frozen DAQ issue, arising from loss of communication with cdaql3. Expert was called (B. Sawatzky), who came to counting house, power cycled cdaql3, ran an fsck and re-booted all ROCs (https://logbooks.jlab.org/entry/3630647) (thank you Brad).
- Day shift continue to take data on Group 8, and will soon move to Group 7 (Q2 = 3.979, x = 0.5, pi+).
- The plan after this is Groups 3 (Q2=4, x=0.4, pi+) and 4 (Q2=4, x=0.4, pi-), followed by 17 (Q2=5.5, x=0.5, z=0.7, pi+ and pi-) and 18 (Q2=5.5, x=0.5, z=0.6, pi+ and pi-).

RC Meeting:
- We have no planned RC meetings until Monday. Unless there is a special request.

Notices in effect for shift crews:
* NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''
* Shift leaders, please fill out the check list once per shift!
* It is no longer required to fill out the standard.kinematics file online, it is now done automatically
* Split the large desired charge between multiple runs (~1 hour per run)
* Listed currents are estimates. For runs at less than 70uA, increase current until SHMS singles rates are between 450k-500k.
* Remember to check trigger pre-scales before runs (PS6=0 all the time, adjusting PS4 to give at least 100Hz, and no more than 200Hz, of HMS singles, PS1=PS2=PS3=PS5=-1).
* Check that the pre-scaler gui is active and hasn't frozen/doesn't need re-started, especially between target changes
* How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Monday, Nov 26: Energy change to 1.86 GeV/pass and move to 500 MHz laser. Hall will go to restricted access Monday morning and aim is to be locked up again Tuesday Swing. There has been a separate e-mail about shift updates. TO needed for all shifts. SL cancelled Mon day 11/26, Mon swing 11/26, Tue owl 11/27, Tue day 11/27. SL on standby Tue swing 11/27, Wed owl 11/28 and Wed day 11/28. SL shifts normal from Wed swing 11/28 currently. If on standby please be ready to come to counting house. We will know more Tue/Wed about expected beam delivery - stay tuned for updates to schedule.

Planned accesses to the Hall:
Monday, Nov 26: Work in Hall should be coordinated with Walter K. Aerogel change by V. Berdnikov (~8 hours duration Mon). HGC work by B. Sawatzky and University of Regina team. D. Meekins will investigate target loop 3 JT valve issue. V. Kumar (Regina group) walkthrough Monday morning.
Tuesday, Nov 27: installation of brushes near the spectrometer wheels for S. Lassiter (originally planned for Tue/Wed).
Please let me know any other plans for work in the Hall Mon/Tue.

Special requests:
Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463). Has to be done before shut down in Dec.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3630770
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181124/918ad7ca/attachment.html>

More information about the Hallc_running mailing list